The perimeter of a tennis table is 28 feet. the difference between 4 times the length and 3 times the width is 14 feet. find the dimensions.

Answer by macston(5194)

W=width; L=length; P=perimeter=28 feet
P=2(L+W)
28ft=2(L+W)
28 feet=2L+2W Multiply each side by 2.
56 feet=4L+4W Swap sides for clarity.
4L+4W=56 feet
4L-3W=14 feet Subtract equations.
7W=42 feet Divide each side by 7.
W=6 feet ANSWER 1: The width is 6 feet.
P=2(L+W)
28ft=2(L+6ft)
28ft=2L+12 ft
16ft=2L
8ft=L ANSWER 2: The length is 8 feet.
CHECK:
4L-3W=14 ft
4(8ft)-3(6ft)=14ft
32ft-18ft=14ft
14ft=14ft